Two cards are drawn without replacement from an ordinary deck of 52 playing cards. What is the probability that both are spades if the first card drawn was a spade?
I got so far that there are 13/52 and then you take away one giving you 12/52 and then I am lost.

Found 3 solutions by stanbon, gerrylucas, edjones:
Answer by stanbon(75887)

Two cards are drawn without replacement from an ordinary deck of 52 playing cards. What is the probability that both are spades if the first card drawn was a spade?
I got so far that there are 13/52 and then you take away one giving you 12/52 and then I am lost.
------------------
Yes, 13/52; but then 12/51.
Now multiply those fractions
-------------------------------
Using formulas you would get:
P(spade and then spade) = P(spade)*P(spade | spade) = (13/52)(12/51)= 0.0588..
===============================
Cheers,
Stan H.

Answer by gerrylucas(2)

39/638 or about 6.11%

Answer by edjones(8007)

Since the first is a spade the probability that the second is a spade is 12/51.
.
Ed